QUESTION 14 Which of the following statements is captured by the accounting identity, S+ (T-G)+CA? A country with a low savings rate and a government budget deficit needs investment from other countries to supplement its domestic investment funds. A country with a high savings rate and a government surplus has additional funds available to invest in other countries A country that would like to increase its government budget deficit and domestic investment or increase private savings. reduce its current account deficit must lower All of the above.
Answer #1

The given accounting identity is as follows -

S + (T-G) = I + CA

This identity can be written as follows -

S + (T-G) - I = CA

This indicates that sum total of private domestic saving and the public saving exceeds the domestic investment.

In such case, excess savings will be invested abroad.

Thus,

The given identity states that a country with a high savings rate and a government surplus has additional funds available to invest in other countries.

Hence, the correct answer is the option (2).